# LLM-Explorer: A Plug-in Reinforcement Learning Policy Exploration Enhancement Driven by Large Language Models

### v1

@@ Abstract
Policy exploration is critical in reinforcement learning (RL), where existing approaches include $\epsilon$-greedy, Gaussian process, etc. However, these approaches utilize preset stochastic processes and are indiscriminately applied in all kinds of RL tasks without considering task-specific features that influence policy exploration. Moreover, during RL training, the evolution of such stochastic processes is rigid, which typically only incorporates a decay in the variance, failing to adjust flexibly according to the agent’s real-time learning status. Inspired by the analyzing and reasoning capability of large language models (LLMs), we design LLM-Explorer to adaptively generate task-specific exploration strategies with LLMs, enhancing the policy exploration in RL. In our design, we sample the learning trajectory of the agent during the RL training in a given task and prompt the LLM to analyze the agent’s current policy learning status and then generate a probability distribution for future policy exploration. Updating the probability distribution periodically, we derive a stochastic process specialized for the particular task and dynamically adjusted to adapt to the learning process. Our design is a plug-in module compatible with various widely applied RL algorithms, including the DQN series, DDPG, TD3, and any possible variants developed based on them. Through extensive experiments on the Atari and MuJoCo benchmarks, we demonstrate LLM-Explorer’s capability to enhance RL policy exploration, achieving an average performance improvement up to 37.27%. Our code is open-source at https://anonymous.4open.science/r/LLM-Explorer-19BE for reproducibility.
@@ 4.1 Experimental Settings
We evaluate the performance of LLM-Explorer on the Atari (bellemare2013arcade, ; kaiser2019model, ) and MuJoCo (todorov2012mujoco, ) benchmarks, covering tasks with both discrete and continuous action spaces. In the main experiments, we use DQN (mnih2015human, ) and DDPG (Lillicrap2015DDPG, ) as the basis on Atari and MuJoCo, respectively, and plug our LLM-Explorer into them. We selected 15 out of 26 tasks from Atari and 5 out of 11 from MuJoCo, where raw DQN or DDPG algorithm can converge stably and obtain good rewards. In addition, we set the number of training steps to 100k-500k across different tasks based on how fast the reward increases when training the original DQN or DDPG algorithm. In the LLM-Explorer module, we use GPT-4o mini [^1] https://platform.openai.com/docs/models/gpt-4o-mini as the core LLM and set the two key parameters in our design, namely action sampling density and exploration adjusting interval, as $M=100$ and $K=1$. In our deployment, we fix a set of hyper-parameters across all environments. For reproducibility, we provide specific values of all hyper-parameters in Appendix A and list detailed contents of the prompts in Appendix G.
@@ 4.2 Overall Performance
In our design, LLM-Explorer is a simple plug-in method that can be seamlessly integrated with a wide range of existing RL algorithms. To verify, besides the basic algorithms aforementioned, we selected another three widely applied variants of DQN (see Appendix D), including Double-Dueling DQN (van2016deep, ; wang2016dueling, ), Rainbow (hessel2018rainbow, ), and CURL (laskin2020curl, ). We also include TD3 (fujimoto2018addressing, ), a commonly used upgrade of DDPG. Respectively from the Atari and MuJoCo tasks, we selected three environments with relatively good training outcomes as representatives. In the selected tasks, we train agents with the original versions of the above RL algorithms, as well as the versions integrating our LLM-Explorer module. In each experiment, we repeat the training process with three different random seeds and average the results. We show the learning curves for the 12 experiments (4 algorithms×$\times$3 environments) in Figure 3. As the results illustrate, different RL algorithms exhibit diverse performance in different environments, while LLM-Explorer consistently improves their performance. This proves LLM-Explorer’s compatibility with various RL algorithms, indicating its potential in tasks with either discrete or continuous action spaces.
@@ 4.3 Compatibility with Different LLMs
In the framework work of LLM-Explorer, we utilize the LLMs with text-only prompts, leveraging their text-processing capability to derive smart policy exploration strategies. Instead of relying on some specific types or versions of LLMs, our design is a general framework that can work with various types of LLMs. To evaluate this, besides GPT-4o mini used above, we test several other LLMs that are most widely known, including GPT-4o [^2] https://platform.openai.com/docs/models/gpt-4o, GPT-3.5333https://platform.openai.com/docs/models, Llama-3.1-405B, and Llama-3.1-70B [^4] https://ai.meta.com/blog/meta-llama-3-1. We train agents with the original DQN algorithm and then integrate DQN with our LLM-Explorer method, where the latter is driven by each of these different LLMs. In each experiment, we repeat the training process with three different random seeds and average the results. We summarize the game scores obtained at the end of training in Table 1 and show the learning curves in these experiments in Appendix E. In the results, our method consistently improves the human-normalized score of the original algorithms despite the type of LLMs, indicating its strong compatibility with different LLMs. We observe that GPT-4o mini tends to be the best choice for LLM-Explorer, while the Llama model may outperform others in specific environments. It is also interesting to note that the performance of LLM-Explorer is much worse when driven by GPT-4o than when driven by GPT-4o mini. The actual reason for this is worth future study, while one possible speculation is that the super LLMs, like GPT-4o, are too sophisticated and tend to greedily fit specific actions instead of providing flexible policy exploration with randomness, thus limiting the performance.
4.4 Performance VS Computational Consumption
@@
There exist two major trade-offs between the performance and computational cost of LLM-Explorer, where the first lies in the design of LLM workflow. To uncover the roles of key components in the LLM workflow, we conduct ablation experiments. In one experiment, we remove the summarize & suggestion mechanism and allow a single LLM to directly output a probability distribution for future policy exploration based on the {TaskDescription}, {ActionSequence}, and {EpisodeReward}. In another experiment, we retain the two-stage design of the LLM workflow but do not provide the {TaskDescription}, only informing the LLMs of the environment’s name. In each experiment, we repeat the training process with three different random seeds and average the results. As shown in Table 2 and Appendix E, both ablations continue to improve the performance of the original DQN algorithm while significantly reducing the token consumption of LLM. However, the first ablation lacks sufficient analysis of the agent’s learning status, making it less flexible for adjustment during the training process. The second ablation lacks sufficient environmental information, making it less adaptive to specific environments. As a result, neither of them performs as well as the full design. We provide a more detailed discussion on the role of task description in the prompt in Appendix F.
@@
The second trade-off lies in the setting of the two key parameters, namely action sampling density ($M$) and exploration adjusting interval ($K$). By reducing sampling density, i.e., smaller $M$, or reducing the frequency of adjusting the exploration strategy, i.e., larger $K$, we can obviously reduce the token consumption of LLM. To evaluate the impact of these, we conduct experiments and show the results in Table 3 and Appendix E. As the results illustrate, LLM-Explorer with either smaller $M$ or larger $K$ keeps improving the performance of the original DQN algorithm. However, smaller $M$ provides insufficient information about the agent’s real-time learning status, and larger $K$ limits adjustments to the exploration strategy. As a result, both of them are less capable of flexibly adapting the policy exploration to the training process, achieving worse performance than LLM-Explorer with the original settings of $M$ and $K$. Moreover, we also analyze the impact of increasing the sampling density, i.e., larger $M$. As the results indicate, although increasing the token consumption of LLM, a larger $M$ does not consistently improve the performance of LLM-Explorer. This may be because the original settings of $M$ already provide sufficient information about the agent’s real-time learning status. Therefore, further increasing the sampling density complicates the LLM’s ability to analyze and summarize the data, which may hinder overall performance.
@@
From these analyses, we demonstrate that the full design and properly configured values of $M$ and $K$ are critical for achieving the best performance of LLM-Explorer. However, we also highlight the trade-offs between performance and computational consumption in LLM-Explorer. Therefore, for deployments with limited computational resources, it is possible to simplify the design of the LLM workflow or adjust $M$ and $K$ as above to reduce computational consumption while still maintaining certain performance improvements over the original RL algorithm. For deployments with sufficient computational resources, the full design with the original settings of $M$ and $K$ is the optimal choice.

### v2

@@ Abstract
Policy exploration is critical in reinforcement learning (RL), where existing approaches include $\epsilon$-greedy, Gaussian process, etc. However, these approaches utilize preset stochastic processes and are indiscriminately applied in all kinds of RL tasks without considering task-specific features that influence policy exploration. Moreover, during RL training, the evolution of such stochastic processes is rigid, which typically only incorporates a decay in the variance, failing to adjust flexibly according to the agent’s real-time learning status. Inspired by the analyzing and reasoning capability of large language models (LLMs), we design LLM-Explorer to adaptively generate task-specific exploration strategies with LLMs, enhancing the policy exploration in RL. In our design, we sample the learning trajectory of the agent during the RL training in a given task and prompt the LLM to analyze the agent’s current policy learning status and then generate a probability distribution for future policy exploration. Updating the probability distribution periodically, we derive a stochastic process specialized for the particular task and dynamically adjusted to adapt to the learning process. Our design is a plug-in module compatible with various widely applied RL algorithms, including the DQN series, DDPG, TD3, and any possible variants developed based on them. Through extensive experiments on the Atari and MuJoCo benchmarks, we demonstrate LLM-Explorer’s capability to enhance RL policy exploration, achieving an average performance improvement up to 37.27%. Our code is open-source at https://github.com/tsinghua-fib-lab/LLM-Explorer for reproducibility.
@@ 4.1 Experimental Settings
We evaluate the performance of LLM-Explorer on the Atari (bellemare2013arcade, ; kaiser2019model, ) and MuJoCo (todorov2012mujoco, ) benchmarks, covering tasks with both discrete and continuous action spaces. In the main experiments, we use DQN (mnih2015human, ) and DDPG (Lillicrap2015DDPG, ) as the basis on Atari and MuJoCo, respectively, and plug our LLM-Explorer into them. We selected 15 out of 26 tasks from Atari and 5 out of 11 from MuJoCo, where raw DQN or DDPG algorithm can converge stably and obtain good rewards. In addition, we set the number of training steps to 100k-500k across different tasks based on how fast the reward increases when training the original DQN or DDPG algorithm. In the LLM-Explorer module, we use GPT-4o mini [^1] https://platform.openai.com/docs/models/gpt-4o-mini as the core LLM and set the two key parameters in our design, namely action sampling density and exploration adjusting interval, as $M=100$ and $K=1$. In our deployment, we fix a set of hyper-parameters across all environments. For reproducibility, we provide specific values of all hyper-parameters in Appendix A and list detailed contents of the prompts in Appendix L.
@@ 4.2 Overall Performance
[…]
4.3 Computational Overhead
Table 1: Wall-clock time and API cost for a 500k-step run across all Atari and MuJoCo tasks. For each task, we average 3 repeated trainings with different random seeds.
Environment Task Input token (k) Output token (k) Cost (\$) Time (h)
Atari Alien 1243.65 897.95 0.73 9.23
Amidar 1563.54 927.00 0.79 9.37
BankHeist 1514.98 1101.00 0.89 9.42
Breakout 3821.06 2634.00 2.15 11.02
ChopperCommand 1340.64 684.00 0.61 9.18
CrazyClimber 413.23 208.00 0.19 8.59
Freeway 366.87 231.25 0.19 8.58
Hero 1191.81 686.00 0.59 9.12
Jamesbond 2133.88 1011.00 0.93 9.64
Krull 805.92 438.00 0.38 8.85
MsPacman 1456.50 1006.10 0.82 9.36
Pong 260.83 312.00 0.23 8.57
Qbert 2263.39 1626.00 1.32 9.95
Seaquest 1389.57 768.00 0.67 9.23
UpNDown 1447.75 1011.00 0.82 9.36
MuJoCo HalfCheetah 978.16 524.98 0.46 8.96
Hopper 5693.25 3611.96 3.02 12.21
Humanoid 25387.59 7477.76 8.29 22.03
HumanoidStandup 2061.12 610.47 0.68 9.45
Walker2d 4422.74 2475.20 2.15 11.21
Average 2987.82 1412.08 1.30 10.17
In this section, we analyze the computational time and cost, which is crucial for assessing the practical application potential of our method. As shown in table 1, on average, each training run consumes approximately \$1.3 in API calls and takes about 10 hours to complete. We consider this to be an acceptable and reasonable overhead. We also provided more detailed computational overhead comparison with baseline methods in Appendix F.
@@ 4.4 Compatibility with Different LLMs
In the framework work of LLM-Explorer, we utilize the LLMs with text-only prompts, leveraging their text-processing capability to derive smart policy exploration strategies. Instead of relying on some specific types or versions of LLMs, our design is a general framework that can work with various types of LLMs. To evaluate this, besides GPT-4o mini used above, we test several other LLMs that are most widely known, including GPT-4o [^2] https://platform.openai.com/docs/models/gpt-4o, GPT-3.5333https://platform.openai.com/docs/models, Llama-3.1-405B, and Llama-3.1-70B [^4] https://ai.meta.com/blog/meta-llama-3-1. We train agents with the original DQN algorithm and then integrate DQN with our LLM-Explorer method, where the latter is driven by each of these different LLMs. In each experiment, we repeat the training process with three different random seeds and average the results. We summarize the game scores obtained at the end of training in Table 2 and show the learning curves in these experiments in Appendix E. In the results, our method consistently improves the human-normalized score of the original algorithms despite the type of LLMs, indicating its strong compatibility with different LLMs.
4.5 Balance between Performance and Consumption
@@
Components in the LLM workflow. To uncover the roles of key components in the LLM workflow, we conduct ablation experiments. In one experiment, we remove the summarize & suggestion mechanism and allow a single LLM to directly output a probability distribution for future policy exploration based on the {TaskDescription}, {ActionSequence}, and {EpisodeReward}. In another experiment, we retain the two-stage design of the LLM workflow but do not provide the {TaskDescription}, only informing the LLMs of the environment’s name. In each experiment, we repeat the training process with three different random seeds and average the results. As shown in Table 3 and Appendix E, both ablations continue to improve the performance of the original DQN algorithm while significantly reducing the token consumption of LLM. However, the first ablation lacks sufficient analysis of the agent’s learning status, making it less flexible for adjustment during the training process. The second ablation lacks sufficient environmental information, making it less adaptive to specific environments. As a result, neither of them performs as well as the full design.
@@
Setting of key parameters. There are two key parameters within the workflow, namely action sampling density ($M$) and exploration adjusting interval ($K$). By reducing sampling density, i.e., smaller $M$, or reducing the frequency of adjusting the exploration strategy, i.e., larger $K$, we can obviously reduce the token consumption of LLM. To evaluate the impact of these, we conduct experiments and show the results in Table 4 and Appendix E. As the results illustrate, LLM-Explorer with either smaller $M$ or larger $K$ keeps improving the performance of the original DQN algorithm. However, smaller $M$ provides insufficient information about the agent’s real-time learning status, and larger $K$ limits adjustments to the exploration strategy. As a result, both of them are less capable of flexibly adapting the policy exploration to the training process, achieving worse performance than LLM-Explorer with the original settings of $M$ and $K$. Moreover, we also analyze the impact of increasing the sampling density, i.e., larger $M$. As the results indicate, although increasing the token consumption of LLM, a larger $M$ does not consistently improve the performance of LLM-Explorer. This may be because the original settings of $M$ already provide sufficient information about the agent’s real-time learning status. Therefore, further increasing the sampling density complicates the LLM’s ability to analyze and summarize the data, which may hinder overall performance.
Table 5: Performance of LLM-Explorer with different LLM sizes. The human-norm scores (%) are recorded at the end of training and averaged across 3 random seeds. The underlines indicate improvements over the raw RL algorithm. The bold fonts are the best results.
Task DQN DQN+LLM-Explorer
Qwen2.5-32B Qwen2.5-14B Qwen2.5-7B Qwen2.5-3B
Alien 0.26 0.51 0.34 0.16 0.06
Freeway 17.75 59.49 30.78 26.32 16.93
MsPacman 1.56 2.60 1.94 1.51 1.14
Size of LLMs. As shown above, the proposed LLM-Explorer framework is compatible with different LLMs. To further explore the impact of LLM size, we test multiple Qwen2.5 models with varying sizes. As shown in Table 5, the performance degrades with smaller models, and models with $\leq$7B parameters generally fail to work. This suggests that, in general, larger and more capable models tend to perform better.
@@
From these analyses, we demonstrate that the full design pipeline, properly configured values of $M$ and $K$, and capable large LLMs are critical for achieving the best performance of LLM-Explorer. However, we also highlight the trade-offs between performance and computational consumption in LLM-Explorer. Therefore, for deployments with limited computational resources, it is possible to simplify the design of the LLM workflow or adjust $M$ and $K$ as above to reduce computational consumption while still maintaining certain performance improvements over the original RL algorithm. For deployments with sufficient computational resources, the full design with the original settings of $M$ and $K$ is the optimal choice. Also, fine-tuning a specialized LLM, for example, distilling the exploration strategy generation from a large LLM into a smaller one, would further reduce the minimum capable LLM size and save more computational cost.
@@ 6 Conclusions
Acknowledgments and Disclosure of Funding
This work was supported in part by the National Key Research and Development Program of China under 2024YFC3307500, and Beijing National Research Center for Information Science and Technology.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The contribution of our paper is clearly summarized in section 1, which reflects our work in the whole paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitation of our work in Appendix M.1.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our paper does not includes theoretical proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We includes the implementation details in Section 4.1 and Appendix A. We also release our codes at https://github.com/tsinghua-fib-lab/LLM-Explorer.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code is available at https://github.com/tsinghua-fib-lab/LLM-Explorer and will be made publicly available after the reviewing procedure.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The details about our training and testing are included in Section 4.1 and Appendix A.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We show standard deviations across multiple random seeds in Section 4.2.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We discussed the computational consumptions in Section 4.3.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We follow the NeurIPS Codes of Ethics, and details are discussed in Appendix M.2.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Our model may have a broader impact, as discussed in Appendix M.3.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: There is no such risks in our paper.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The original paper of models and benchmarks are correctly cited in Section 4.1.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our code is open sourced in  https://github.com/tsinghua-fib-lab/LLM-Explorer.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work does not involve crowdsourcing.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work does not involve crowdsourcing.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs are our research objects, but are not involved in designing and creating the work.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.